OpenStudy (anonymous):

two-thirds times 5 and one-half 2 and two-thirds 3 and two-thirds 4 and one-third 5 and one-third

OpenStudy (anonymous):

\[\frac{2}{3} \times 5 \frac{1}{2} = \frac{2}{3} \times \frac{11}{2}\] Clearly, this cancels to 11/3, which is equivalent to 3 and two thirds.

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

oh, i see

OpenStudy (anonymous):

thnx :D

OpenStudy (anonymous):

here's a medal 4 u